OpenStudy (anonymous):

what percent of 180 is 54

OpenStudy (anonymous):

x(180) = 54 x = 54/180 =.3 = 30%

OpenStudy (anonymous):

54/180 = x / 100 cross multiply and solve for x

OpenStudy (anonymous):

yes i got .3 as well

OpenStudy (anonymous):

which gave me 30%